OpenStudy (anonymous):

solve the equation using the addition property of equality p-1=4

OpenStudy (carson889):

A simple do to one side what you do to the other. Thus to isolate p, we need to get the -1 off the left hand side. So to negate the -1 we must add a positive one to the left side ( and as we know also the right hand side). Therefore, p - 1 + 1 = 4 + 1 Do the addition and subtration. -1 + 1 = 0, 4 + 1 =5. Thus p + 0 = 5, which equals p=5